OpenStudy (anonymous):

can someone please chat with me or something and help me with my quiz? it's 15 questions and I have 13 to go. The quiz is on solving linear equations with alegebra

OpenStudy (anonymous):

sure what is is

OpenStudy (anonymous):

what is the resulting equation when the expression for y in the second equation is substituted into the first equation? 3x+y=1 y=6-4x A. -x+6=1 B. x+6=1 C. 7x+6=1

OpenStudy (whpalmer4):

The second equation is \(y = 6-4x\). Wherever you see \(y\) in \(3x+y=1\), you replace it with \((6-4x)\). What is the result?

OpenStudy (anonymous):

is it 7x+6=1?

OpenStudy (whpalmer4):

No. \[3x+y=1\]Replace \(y\) with \((6-4x)\): \[3x + (6-4x) = 1\]\[3x+6 - 4x = 1\]Collect like terms and you get...

OpenStudy (anonymous):

-x+6=1? because you add the 3x and -4x?

OpenStudy (anonymous):

omgomgomgomgomggggg i'm gonna dieeeeee.

OpenStudy (whpalmer4):

yes, -x+6 =1 is correct!